The vapour pressure of pure liquid A is 10 torr and at the same temperature when 1g of B solid is dissolved in 20g of A,its vapour pressure is reduced to 9 torr if the molecular mass of A is 200amu then the molecular mass of B is 

Grade:12

1 Answers

Arun
6 years ago
Dear Student
 
x2 = – \Deltap2/ pi = 1 / 10 = 0.1
now 
x2 = [(m2 /M2) /(m1/ M1)]
0.1 = [(1/M2) / (20/200)]
M2 = 100 g mol-1
 
Regards
Arun (askIITians forum expert)